# Reframing Image Difference Captioning with BLIP2IDC and Synthetic Augmentation

### v1

@@ 4 Datasets
The structure of IDC datasets is a triplet $(I_{ref},I_{modified},{GT})$ which respectively represents the original image $I_{ref}$, the modified image $I_{modified}$, and the ground-truth set of reference descriptions of the differences $GT$. We compare our approach to recent IDC models on standard IDC datasets: CLEVR-Change [29], CLEVR-DC [13], Spot-The-Diff [31] and Image Editing Request [41]. These datasets are distinguished by various properties, summarized in Tab. 1: the number of image pairs ($I_{ref},I_{modified}$), the origin of the $I_{ref}$ images, which may be real images or generated by 3D models, the number of reference captions (ground-truths $GT$) per triplet, the way in which $I_{modified}$ and $GT$ are obtained, the number and type of transformations performed, and whether some manual intervention was required to create the dataset. This may happen at the level of triplet filtering or the creation of ground-truth legends.
These properties can limit the creation of these datasets, which can be costly, as well as their usefulness, for example when the diversity of transformation types is reduced. We discuss the limitations of these datasets hereafter, as well as the issues posed by the evaluation of the IDC task. To handle those limitations, we propose a method based on text-guided image editing methods for creating IDC datasets without manual intervention, which can then be adapted to different use cases. Furthermore, instruction-based image editing capabilities are evaluated on some benchmarks (IE datasets), like MagicBrush [53] or Emu Edit test set generations (EE) [37]. Those datasets are also composed of triplets $(I_{ref},T_{instruction},I_{edited})$, with $T_{instruction}$ the modification that maps $I_{ref}$ to $I_{edited}$ and $I_{edited}$ the edited image. We thus propose to adapt those datasets for the IDC task.
@@
EE [37] 5,612 Real-world Automatic generation $\checkmark$ 1.0
@@ 4.1 Existing datasets
Emu Edit (EE). This dataset [37] from Meta is composed of pairs of original images from MSCOCO [20] and modified images generated by the Emu Edit model, a state-of-the-art generation model, which takes an image and an editing instruction as inputs. This dataset also contains fine-grained information about the type of modifications generated, grouped in eight different categories : Add, Text, Color, Background, Local Style, Global and Remove. Paired with the real-world setting of original images which enables for greater usability, these properties make it a great candidate for a challenging IDC benchmark.
@@ 4.2 IDC task issues
IDC models are trained on triplets $(I_{ref},I_{modified},GT)$ and evaluated on a test set of pairs $(I_{ref},I_{modified})$ by comparing the generated caption with the reference captions $GT$, using various automatic evaluation metrics, such as BLEU, ROUGE, METEOR and CIDEr [28, 19, 15, 27]. Some aspects of the real-world datasets weaken the quality of the evaluation.
@@ 4.3 Synthetic augmentation
To evaluate this pipeline, we applied it to augment the EE dataset [37], as this dataset provides real-world edited images alongside editing instructions well-suited for an image editing model and with a clear distinction between each class of modifications. As image editing model, we choose the InstructPix2Pix [2] model fine-tuned on the MagicBrush [53] dataset, the state-of-the-art solution for executing real-world edits. The train set is augmented by generating eight new modified images per original ones from the EE dataset and their associated editing instructions, without any manual curation.
To ensure meaningful evaluation, we leverage the Llama-2-7b-chat-hf model to generate 4 additional variations of each editing instructions from the EE test set (cf. supplementary). Each of the 2,022 samples in the test set is therefore described by 5 GT captions: the original instruction and the 4 reference captions provided by the LLM. Note that the train set images (see Fig. 5) were not modified by the same editing model as in the test set.
Figure 5: Comparison of train set samples from EE, generated by the Emu Edit model, and Syned, generated by a fine-tuned InstructPix2Pix model.
@@ 5.1 Evaluation protocol
Models and datasets. We compare BLIP2IDC to other state-of-the-art IDC models with the results reported in their respective paper on the four standard IDC datasets: CLEVR-Change, CLEVR-DC, STD, and IER. We also introduce the EE dataset to the IDC task, on which we also train and evaluate CLIP4IDC [6] and SCORER [44] for the sake of comparison.
@@
BLIP2IDC is fine-tuned as described in Sec. 3.2, with a LoRA of rank $8$. For our own adaptation of BLIP2, we use the decoder-only BLIP2’s LLM version, with opt2.7B and vit-base-patch16-224 as the ViT model. We use LoRA implementation from the peft[24] python library to train BLIP2 and to store adaptation weights in a lightweight 20 MB file. For CLIP4IDC and SCORER, we either used the hyper-parameters from their respective online repository, or hyperparameters communicated by the authors (see supplementary).
@@ 5.3 Automatically generated dataset
We introduced in Sec. 4.3 Syned, a new dataset designed for the IDC task, built upon the EE dataset. New data are generated by a different instruction-based image editing model, thus broadening the scope and diversity of the new modified images wrt the original EE dataset, as depicted in Fig. 5. We evaluate the usefulness of Syned as data augmentation for different IDC models in Tab. 4. We show an improvement with the proposed synthetic augmentation (EE + Syned) for both CLIP4IDC and BLIP2IDC. Note that the available implementation of SCORER enables to reach at most a CIDEr score of 23.2 which is not significant in this setting and can be achieved only by reward hacking.
Table 4: CIDEr scores on Emu Edit test set. We show consistent improvements with synthetic augmentation (EE + Syned) on all models.
Models EE Syned+ EE Improvement (%)
@@
EE 101.01 147.69 119.48 155.76 30.53 29.46 93.39 76.24 100.83
Syned+ EE 107.38 147.42 112.37 164.40 37.07 30.43 111.16 77.81 106.83
In Table 5, we study BLIP2IDC behaviour with respect to its training data, being either Syned, EE or both on different modification categories. Combining datasets is generally the best configuration overall, except for the text (T) category. We assume this is due to the poor capacity of most Image Editing models to generate text in images, which hampers the performance of the synthetic augmentation. The worst categories in terms of CIDEr score performance are the Style and Global categories, which are very subjective changes. More details are available in the supplementary material.
@@ 5.4 Qualitative analysis
We present in Fig. 1 a comparative analysis of state-of-the-art IDC models outputs across different scenarii: in-distribution image-pairs from the train and test set of EE, and out-of-distribution images from the web, to assess the zero-shot capabilities of the models.
We observe SCORER’s limitations in handling real-world images in all cases. It only succeed in identifying a text modification in the train sample, and confuses actions like adding or removing. CLIP4IDC performs well on samples from EE but struggles with zero-shot generalization. Specifically, it misidentifies the ”table tennis set” due to its absence in the training data, confusing it with a similar but incorrect object. Additionnaly, it fails to understand the context in which the object is added. BLIP2IDC, on the other hand, excels on data in-distribution and demonstrates superior capability in generalization through effective use of cross-attention. It accurately recognizes and describes the entire scenes in Fig. 1(c-d), including complex additions like a table tennis set, by leveraging its pretraining. This allows BLIP2IDC to not only identify the presence of fire in Fig. 1(d) but also specify the object affected, showcasing its ability to connect semantic changes with their context.

### v2

@@ 4 Datasets
The structure of IDC datasets is a triplet $(I_{ref},I_{modified},{GT})$ which respectively represents the original image $I_{ref}$, the modified image $I_{modified}$, and the ground-truth set of reference descriptions of the differences GTGT. We compare our approach to recent IDC models on standard IDC datasets: CLEVR-Change [29], CLEVR-DC [13], Spot-The-Diff [31] and Image Editing Request [41]. These datasets are distinguished by various properties, summarized in Tab. 1: the number of image pairs ($I_{ref},I_{modified}$), the origin of the $I_{ref}$ images, which may be real images or generated by 3D models, the number of reference captions (ground-truths GTGT) per triplet, the way in which $I_{modified}$ and GTGT are obtained, the number and type of transformations performed, and whether some manual intervention was required to create the dataset. This may happen at the level of triplet filtering or the creation of ground-truth legends.
These properties can limit the creation of these datasets, which can be costly, as well as their usefulness, for example when the diversity of transformation types is reduced. We discuss the limitations of these datasets hereafter, as well as the issues posed by the evaluation of the IDC task. To handle those limitations, we propose a method based on text-guided image editing methods for creating IDC datasets without manual intervention, which can then be adapted to different use cases. Furthermore, instruction-based image editing capabilities are evaluated on some benchmarks (IE datasets), like MagicBrush [53] or Emu Edit test set generations ($E$) [37]. Those datasets are also composed of triplets $(I_{ref},T_{instruction},I_{edited})$, with $T_{instruction}$ the modification that maps $I_{ref}$ to $I_{edited}$ and $I_{edited}$ the edited image. We thus propose to adapt those datasets for the IDC task.
@@
$E$ [37] 5,612 Real-world Automatic generation $\checkmark$ 1.0
@@ 4.1 Existing datasets
Emu Edit ($E$). This dataset [37] from Meta is composed of pairs of original images from MSCOCO [20] and modified images generated by the Emu Edit model, a state-of-the-art generation model, which takes an image and an editing instruction as inputs. This dataset also contains fine-grained information about the type of modifications generated, grouped in eight different categories : Add, Text, Color, Background, Local Style, Global and Remove. Paired with the real-world setting of original images which enables for greater usability, these properties make it a great candidate for a challenging IDC benchmark.
@@ 4.2 IDC task issues
IDC models are trained on triplets $(I_{ref},I_{modified},GT)$ and evaluated on a test set of pairs $(I_{ref},I_{modified})$ by comparing the generated caption with the reference captions GTGT, using various automatic evaluation metrics, such as BLEU, ROUGE, METEOR and CIDEr [28, 19, 15, 27]. Some aspects of the real-world datasets weaken the quality of the evaluation.
@@ 4.3 Synthetic augmentation
To evaluate this pipeline, we applied it to augment the $E$ dataset [37], as this dataset provides real-world edited images alongside editing instructions well-suited for an image editing model and with a clear distinction between each class of modifications. As image editing model, we choose the InstructPix2Pix [2] model fine-tuned on the MagicBrush [53] dataset, the state-of-the-art solution for executing real-world edits. The train set is augmented by generating eight new modified images per original ones from the $E$ dataset and their associated editing instructions, without any manual curation.
To ensure meaningful evaluation, we leverage the Llama-2-7b-chat-hf model to generate 4 additional variations of each editing instructions from the $E$ test set (cf. supplementary). Each of the 2,022 samples in the test set is therefore described by 5 GT captions: the original instruction and the 4 reference captions provided by the LLM. Note that the train set images (see Fig. 5) were not modified by the same editing model as in the test set.
Figure 5: Comparison of train set samples from $E$, generated by the Emu Edit model, and Syned, generated by a fine-tuned InstructPix2Pix model.
@@ 5.1 Evaluation protocol
Models and datasets. We compare BLIP2IDC to other state-of-the-art IDC models with the results reported in their respective paper on the four standard IDC datasets: CLEVR-Change, CLEVR-DC, STD, and IER. We also introduce the $E$ dataset to the IDC task, on which we also train and evaluate CLIP4IDC [6] and SCORER [44] for the sake of comparison.
@@
BLIP2IDC is fine-tuned as described in Sec. 3.2, with a LoRA of rank 88. For our own adaptation of BLIP2, we use the decoder-only BLIP2’s LLM version, with opt2.7B and vit-base-patch16-224 as the ViT model. We use LoRA implementation from the peft[24] python library to train BLIP2 and to store adaptation weights in a lightweight 20 MB file. For CLIP4IDC and SCORER, we either used the hyper-parameters from their respective online repository, or hyperparameters communicated by the authors (see supplementary).
@@ 5.3 Automatically generated dataset
We introduced in Sec. 4.3 Syned, a new dataset designed for the IDC task, built upon the $E$ dataset. New data are generated by a different instruction-based image editing model, thus broadening the scope and diversity of the new modified images wrt the original $E$ dataset, as depicted in Fig. 5. We evaluate the usefulness of Syned as data augmentation for different IDC models in Tab. 4. We show an improvement with the proposed synthetic augmentation ($E$ + Syned) for both CLIP4IDC and BLIP2IDC. Note that the available implementation of SCORER enables to reach at most a CIDEr score of 23.2 which is not significant in this setting and can be achieved only by reward hacking.
Table 4: CIDEr scores on Emu Edit test set. We show consistent improvements with synthetic augmentation ($E$ + Syned) on all models.
Models $E$ Syned+ $E$ Improvement (%)
@@
$E$ 101.01 147.69 119.48 155.76 30.53 29.46 93.39 76.24 100.83
Syned+ $E$ 107.38 147.42 112.37 164.40 37.07 30.43 111.16 77.81 106.83
In Table 5, we study BLIP2IDC behaviour with respect to its training data, being either Syned, $E$ or both on different modification categories. Combining datasets is generally the best configuration overall, except for the text (T) category. We assume this is due to the poor capacity of most Image Editing models to generate text in images, which hampers the performance of the synthetic augmentation. The worst categories in terms of CIDEr score performance are the Style and Global categories, which are very subjective changes. More details are available in the supplementary material.
@@ 5.4 Qualitative analysis
We present in Fig. 1 a comparative analysis of state-of-the-art IDC models outputs across different scenarii: in-distribution image-pairs from the train and test set of $E$, and out-of-distribution images from the web, to assess the zero-shot capabilities of the models.
We observe SCORER’s limitations in handling real-world images in all cases. It only succeed in identifying a text modification in the train sample, and confuses actions like adding or removing. CLIP4IDC performs well on samples from $E$ but struggles with zero-shot generalization. Specifically, it misidentifies the ”table tennis set” due to its absence in the training data, confusing it with a similar but incorrect object. Additionnaly, it fails to understand the context in which the object is added. BLIP2IDC, on the other hand, excels on data in-distribution and demonstrates superior capability in generalization through effective use of cross-attention. It accurately recognizes and describes the entire scenes in Fig. 1(c-d), including complex additions like a table tennis set, by leveraging its pretraining. This allows BLIP2IDC to not only identify the presence of fire in Fig. 1(d) but also specify the object affected, showcasing its ability to connect semantic changes with their context.